OpenStudy (el_arrow):

use the alternating series test (if possible) to determine whether the series converges or diverges.

OpenStudy (el_arrow):

\[\sum_{n=1}^{\infty} \frac{ (-1)^n }{ \ln(n+12) }\]

OpenStudy (el_arrow):

@rational could you help me with a few more this time they are alternating series test

OpenStudy (el_arrow):

like i know an = 1/ln(n+12) and you get rid of the (-1)^n

OpenStudy (el_arrow):

its equal to zero so it converges right?

OpenStudy (rational):

Yes but you need to show that it passes the requirements for applying alternating series test

OpenStudy (rational):

http://gyazo.com/8ca143e34a12e6b8f6fa658c9ff877e7

OpenStudy (rational):

1. \[\lim\limits_{n\to\infty} \dfrac{1}{\ln(n+2)} = 0\] so requirement 1 is met. 2. show that the sequence \(\left\{\frac{1}{\ln(n+2)}\right\}\) is decreasing

OpenStudy (el_arrow):

\[\frac{ 1 }{ \ln(n+1+12) } < \frac{ 1 }{ \ln(n+12) }\]

OpenStudy (el_arrow):

is that how i show it?

OpenStudy (rational):

that will do!

OpenStudy (rational):

\[\ln(n+12+1) \gt \ln(n+12)~~ \implies~~ \frac{1}{\ln(n+12+1) } \lt \frac{1}{\ln(n+12)}\] so the sequence is decreasing

OpenStudy (rational):

we're good to use alternating series test as we have shown that the given series meets the hypothesis

OpenStudy (rational):

Just conclude by saying that the series converges by alternating series test.

OpenStudy (el_arrow):

okay do you mind if we do two more examples?

OpenStudy (rational):

Okay post

OpenStudy (el_arrow):

\[\sum_{n=1}^{\infty} (-1)^n \frac{ 5n^7 + 4 }{ 4n^6 + 3n + 5}\]

OpenStudy (el_arrow):

for this one i got an = \[\frac{ 5n^7 + 4 }{ 4n^6 + 3n +5}\]

OpenStudy (el_arrow):

and what do you do when the degree of the numerator is greater than the denominator

OpenStudy (rational):

when the degree of numerator is greater than the denominator, the graph wont settle... meaning, the terms keep growing/decreasing without any bound. so the limit is infinity

OpenStudy (el_arrow):

so if its infinity than checking if it decreases is not necessary?

OpenStudy (rational):

The first requirement failed, so there is no point in testing second requirement. we cannot use alternating series test, need to try some other test

OpenStudy (el_arrow):

what other test could we use?

OpenStudy (el_arrow):

could we use the limit test?

OpenStudy (rational):

yes try it

OpenStudy (el_arrow):

alright

OpenStudy (el_arrow):

\[\lim_{n \rightarrow \infty} \frac{ 5n^7+4 }{ 4n^6 + 3n +5 }----->\lim_{n \rightarrow \infty} \frac{ 35n^6 }{ 24n^5 + 3 }-------->\frac{ 210n^5 }{120n^4 }\]

OpenStudy (rational):

what happened to (-1)^n

OpenStudy (el_arrow):

i thought we didnt need it anymore since its not part of an

OpenStudy (rational):

in "limit test" you need to take the entire expression.

OpenStudy (rational):

thats not required only in alternatign series test

OpenStudy (el_arrow):

so it would be a negative value \[-\frac{ 35n^6 }{ 24n^5 + 3}\]

OpenStudy (el_arrow):

right?

OpenStudy (rational):

how do you know whether (-1)^n is negative or positive ?

OpenStudy (el_arrow):

its not its indeterminate

OpenStudy (rational):

Exactly!

OpenStudy (rational):

look at it like this : \[\lim_{n \rightarrow \infty} (-1)^n \frac{ 5n^7+4 }{ 4n^6 + 3n +5 } ~~=~~\left(\lim_{n \rightarrow \infty} (-1)^n\right) \cdot \left(\lim_{n \rightarrow \infty} \frac{ 5n^7+4 }{ 4n^6 + 3n +5 } \right)\]

OpenStudy (rational):

you have worked earlier... the second limit is infinite, right ?

OpenStudy (el_arrow):

right

OpenStudy (rational):

and the first limit seems to be toggling between +1 and -1

OpenStudy (rational):

so clearly the overall limit also toggles bbetween +infty and -infty, yes ?

OpenStudy (rational):

whatever it is, we're convinced it certainly doesn't converge yes ?

OpenStudy (el_arrow):

yes

OpenStudy (rational):

So can we say the series diverges by "limit test"

OpenStudy (el_arrow):

yes we can but is it okay if i use another test when its asking to use the alternating series test

OpenStudy (el_arrow):

like do you think my teacher would give me credit if i use another test besides the one he ask to do

OpenStudy (rational):

show ur work for alternating series test, show that the first requirement of alternating series test is not met

OpenStudy (rational):

then mention that you need to use another test as the hypothesis of "alternating series test" is not met

OpenStudy (rational):

you need to show that all in work of this problem

OpenStudy (el_arrow):

oh yes of course

OpenStudy (rational):

go through Example2 problem quick http://tutorial.math.lamar.edu/Classes/CalcII/AlternatingSeries.aspx

OpenStudy (el_arrow):

oh i see similar to mine that problem is

OpenStudy (el_arrow):

so this other one is \[\sum_{n=1}^{\infty} \frac{ (-1)^n * [\sin (3n)]^2}{ n }\]

OpenStudy (el_arrow):

i got an= \[\frac{ [\sin(3n)]^2 }{ n }\]

OpenStudy (rational):

Yes test the first condition http://gyazo.com/8ca143e34a12e6b8f6fa658c9ff877e7

OpenStudy (el_arrow):

\[\lim_{n \rightarrow \infty} \frac{ [\sin(3( \infty ))]^2 }{ \infty } = 0\]

OpenStudy (rational):

what about second condition

OpenStudy (el_arrow):

\[\frac{ [\sin(3(n+1))]^2 }{ n+1 } < \frac{ [\sin(3n)]^2 }{ n }\]

OpenStudy (el_arrow):

it is not true i plug in 2 for n and i got a higher number on the left side

OpenStudy (rational):

yeah strictly speaking, it fails the second condition

OpenStudy (el_arrow):

so it is divergent?

OpenStudy (rational):

lets look at its graph maybe

OpenStudy (el_arrow):

alright

OpenStudy (rational):

Keep in mind, alternating series test is used only to show convergence. failure of conditions in alternating series test doesn't imply divergence.

OpenStudy (rational):